OpenStudy (anonymous):

write the leibniz notation for the expression: The derivative with respect of beta of cos (beta)

OpenStudy (anonymous):

\[d \cos (\beta)/d \beta=-\sin \beta\]

OpenStudy (anonymous):

Leibniz notation is simply the \[d/dx\] formulation of the derivative problem, as opposed to f'(x) notation.